# Haojun Chen    Minghao Liu    Xiaojian Ma    Zailin Ma    Huimin Wu    Chengdong Ma    Yuanpei Chen    Yifan Zhong    Mingzhi Wang    Qing Li    Yaodong Yang

### v1

@@ 4.2 Simulation Environments and datasets
4.3 Falcon is a Plug & Play Acceleration Algorithm.
@@
In the first experiment, we examine our algorithm on the robomimic benchmarks. We present the success rates, NFE, and speedup for both the original diffusion models and their Falcon-enhanced counterparts in Table 1, Table 2, and Table 3, respectively. To better understand Falcon’s impact, we categorize the tasks into simpler and more complex ones based on the level of fine-grained manipulation required. In the Robomimic benchmark (Mandlekar et al., 2021), Lift and Can are relatively simple tasks, while the remaining tasks, Square, Transport, and ToolHang, are more complex.
@@
Simple tasks.  The Lift and Can tasks involve smooth, predictable motions. In Lift, the robotic arm moves downward to grasp a small cube and then upward to lift it, with minimal variation in action within each phase. Can follows a similar pattern, where the robot picks up a coke can from a bin and places it in a target bin, primarily involving steady translational motion. Both tasks exhibit strong dependencies between consecutive actions, making them ideal for Falcon’s acceleration. As shown in Fig.5 (left), Falcon starts denoising with low noise, significantly speeding up the process. Table3 shows Falcon achieving a 7x speedup in Lift and 2-3x in Can, demonstrating its efficiency in structured motion tasks.
@@
Difficult tasks.  Square, Transport, and ToolHang are more complex and involve finer, more precise movements. For example, in the Transport task, two robot arms must work together to transfer a hammer from a closed container on a shelf into a target bin on another shelf. One arm retrieves the hammer from the container, while the other arm first clears the target bin by removing trash. Then, one arm hands over the hammer to the other, which must place it in the target bin. These actions are dynamic and entail larger changes between consecutive movements. As a result, Falcon can’t start denoising actions with low noise level (see Fig. 5 (right)). Consequently, Falcon’s speedup is not as pronounced as in simpler tasks. However, even in these more challenging scenarios, as shown in Table 3, Falcon still achieves around 2x speedup while maintaining performance close to that of the original models.
@@
BlockPush Kitchen
p1 p2 p1 p2 p3 p4
DDPM 0.99 0.94 1.00 0.99 0.99 0.96
DDPM+Falcon 0.99 0.97 1.00 0.99 0.99 0.96
[…]
BlockPush Kitchen
NFE Speedup NFE Speedup
DDPM 100 1.0x 100 1.0x
DDPM+Falcon 32.7 2.8x 19.03 5.25x
4.4 Falcon achieve acceleration diffusion policy in long horizon tasks
@@ 4.5 Falcon retains the ability to express multimodality policy
We investigate whether Falcon might compromise the ability to express multimodal policy similar to the Consistency Policy (Prasad et al., 2024). We conducted experiments in the BlockPush environment, performing 200 evaluation episodes. During these evaluations, we computed the frequency of each block being pushed to each square. Fig. 4 shows that the frequency of different policy modalities is nearly the same, indicating that Falcon does not hinder the policy’s ability to model multimodal action distributions. This uniform distribution suggests that Falcon preserves the multimodality of the diffusion policy, effectively handling different possible action combinations without bias. We also test the performance and speed in this environment. As shown in Table 5 and Table 5, Falcon achieves a 2.8x speedup in this environment. Therefore, we conclude that Falcon maintains the ability to model multimodal policies while accelerating the denoising process, ensuring that the policy can still explore a variety of action sequences.
Figure 3: Left: Ablation analysis on threshold $\epsilon$. The effect of varying the $\epsilon$ threshold on both the score (red circles) and speedup (blue squares), demonstrating the trade-off between performance and speed as $\epsilon$ changes. Middle: Ablation analysis on exploration rate $\delta$. The score (red circles) increases sharply with $\delta$ and then stabilizes, while the speedup (blue squares) decreases as the exploration rate rises. This study highlights the balance between exploration and exploitation. Right: Ablation analysis on selection mechanism. This figure compares Falcon’s performance with different partial denoised action selection configurations. Restricting Falcon to choose actions with fixed noise levels (K/2, K/5) significantly reduces score, highlighting that the selection mechanism ensures accurate action choices while accelerating denoising and maintaining high performance.
@@ 4.6 Ablation analysis: sensitivity to the threshold $\epsilon$
We explore how the threshold $\epsilon$ influences Falcon’s performance in the Square_ph task. The threshold $\epsilon$ determines the confidence in using partial denoised actions from previous steps to the denoising process towards the desired action $\boldsymbol{a}_{t+T_{p}}$ guided by $\nabla_{\mathbf{A}_{t}^{k}}\log p\left(\mathbf{A}_{t}^{k}\mid\mathbf{O}_{t}\right)$. We conducted 50 evaluations using a range of $\epsilon$ values: $\{10^{-4},5\times 10^{-3},8\times 10^{-3},10^{-2},3\times 10^{-2},5\times 10^{-2},8\times 10^{-2},10^{-1}\}$. As shown in Fig. 3 (left), when $\epsilon$ is small, Falcon is highly selective about using partial denoised actions, leading to slow denoising because Falcon frequently samples from the standard normal distribution. This results in no significant improvement. When $\epsilon$ is large, Falcon becomes overly confident in the partial denoised actions, leading to selecting many incorrect actions that do not align with $\boldsymbol{a}_{t+T_{p}}$, which causes a sharp drop in Score. The optimal range of $\epsilon$ lies in the middle, where Falcon can achieve speedup without sacrificing task performance, demonstrating that a balanced $\epsilon$ allows Falcon to maintain both efficiency and accuracy.
4.7 Ablation analysis: sensitivity to exploratian rate $\delta$
We investigate how different exploration rates $\delta$ influence Falcon’s performance by running 50 evaluations across $\{0.001,0.05,0.0625,0.076,0.1,0.2,0.33,1\}$ on Transport ph tasks. Exploration Rate $\delta$ controls the probability that Falcon starts the denoising process from a normal distribution rather than using the partial denoised action from the previous step. This exploration mechanism allows Falcon to fill the latent buffer with more partial denoised actions, increasing exploration and preventing over-reliance on the reference action. As shown in Fig. 3 (middle), when $\delta$ is set to 0, Falcon primarily aligns with the reference action at each step. Since the reference action is not perfectly aligned with the desired action $\mathbf{A}_{t}$, errors accumulate over time, causing Falcon to generate incorrect actions, which results in a significant drop in Score. As $\delta$ increases, the exploration rate encourages Falcon to sample more diverse actions, improving Score by reducing the accumulation of errors. However, further increasing $\delta$ results in a decrease in Speed, as the model becomes less reliant on the history of denoised actions and starts exploring more random samples from the normal distribution.
@@ 4.8 Ablation analysis: sensitivity to selection mechanism
We evaluate Falcon’s partial denoised action selection mechanism. Falcon adaptively selects partial denoised actions based on $\epsilon$ and the one-step estimation, estimating whether each action can be denoised to $A_{t}$. We compare this with the case where Falcon is fixed to always sample from actions with noise levels of $K/2$ or $K/5$ (represented by the red and green bars), without considering the adaptive selection mechanism. Fig. 3 (right) shows a significant decrease in Score when Falcon is restricted to these fixed actions, demonstrating that not leveraging the selection mechanism compromises performance. This confirms that Falcon’s Selection Mechanism is crucial for ensuring accurate action choices, allowing the model to maintain high performance while accelerating the denoising process.
@@ 5 Related Work
Diffusion policies are widely used for modeling complex behaviors in robotics. 3D Diffusion Policy (Ze et al., 2024b) improves visuomotor performance by integrating 3D visual data from sparse point clouds. Large-scale models like Octo (Team et al., 2024) and RDT (Liu et al., 2024) scale diffusion policy parameters to build vision-language-action foundation models. However, their high computational cost limits real-time applications, necessitating acceleration techniques.
To speed up inference, sequential solvers like DDIM (Song et al., 2020a) and DPMSolver (Lu et al., 2022) reduce denoising steps, while ParaDiGMS (Shih et al., 2024) leverages parallel computing to accelerate sampling on GPUs. Distillation-based methods further improve efficiency by training models for single-step inference. Consistency Policy (Kim et al., 2023) aligns diffusion trajectories across timesteps, and ManiCM (Mu et al., 2021) extends this approach to 3D robotic tasks. OneDP (Wang et al., 2024) and SDM (Jia et al., 2024) introduce score-based distillation to reduce performance degradation. However, these methods require task-specific training, limiting adaptability. Streaming Diffusion Policy (SDP) (Høeg et al., 2024) uses partial denoised action trajectories, similar to Falcon. However, SDP requires task-specific training and lacks compatibility with general acceleration techniques.
@@ 6 Conclusion
This paper introduces Falcon, the first diffusion policy approach to accelerate action generation in complex visuomotor tasks by leveraging inter-step dependencies in decision-making. Empirical results confirm that Falcon outperforms strong baselines, such as DDIM and DPMSolver, in terms of speed without sacrificing accuracy or expressiveness. Overall, Falcon offers a simple yet effective solution for real-time robotic tasks, providing efficient action generation for complex visuomotor environments. However, one limitation of Falcon is that it does not use a single set of parameters to accelerate all tasks, since different environments may require task-specific parameter tuning. Future work will explore adaptive parameter selection strategies to further enhance its generalizability across diverse robotic applications.
@@
Appendix A Pseudocode of Falcon with other sampling solvers
In this section, we provide the pseudocode for Falcon when integrated with alternative sampling solvers, such as DDIM (Song et al., 2020a) and DPMSolver (Lu et al., 2022).
A.1 Falcon with DDIM
@@
A.2 Falcon with DPMSolver
@@ Appendix C Detail performance of Falcon with 3D Diffusion Policy
Tables 6, 7 and 8 present the success rates, NFE and speedup respectively, for both 3D Diffusion Policy and 3D FalconDDIM. These results further validate Falcon’s effectiveness in reducing inference time while maintaining task performance across different robotic manipulation tasks.

### v2

@@ 4.2 Simulation and Real-World Environments
Figure 5: Real-world Dexterous Grasping with DDPM+Falcon. Top row (red): head camera view (RealSense D435). Bottom row (green): wrist camera view (RealSense D405C). The sequence shows the successful execution of grasping using Falcon-accelerated policy.
Method Speedup NFE Run Time (s) GPU MEMS (MB) CPU MEMS (MB) Success rate
DDPM $1.0$x $100.00\pm 0.00$ $0.43\pm 0.01$ $3730.17\pm 0.56$ $3737.06\pm 2.63$ 100%
SDP(DDPM) $1.86$x $52.16\pm 2.18$ $0.23\pm 0.01$ $3733.11\pm 2.68$ $3808.67\pm 7.68$ 100%
Falcon+DDPM $\mathbf{3.07x}$ $30.98\pm 1.33$ $0.14\pm 0.01$ $3730.19\pm 0.56$ $3746.67\pm 3.03$ 100%
Table 5: Dexterous Grasping in Real-world Environment. Each entry is evaluated with 20 rollouts in the mean $\pm$ standard deviation format. DDPM and SDP(DDPM) are trained with 1000 epochs with 37 collected expert demonstrations. Our results show that Falcon outperforms SDP in runtime reduction with lower memory usage.
Figure 6: Real-world Square Stick Insertion with DDPM+Falcon. Top row (red): head camera view (RealSense D435). Bottom row (green): wrist camera view (RealSense D405C). The sequence shows the successful execution of insertion using Falcon-accelerated policy.
Method Speedup NFE Run Time (s) GPU MEMS (MB) Success rate
DDPM $1.0$x $100.00_{\pm 0.00}$ $0.43_{\pm 0.01}$ $3735.76_{\pm 0.48}$ 90%
SDP(DDPM) $1.95$x $50.00_{\pm 0.00}$ $0.22_{\pm 0.01}$ $3743.28_{\pm 1.26}$ 85%
Falcon+DDPM $\mathbf{2.86x}$ $25.57_{\pm 7.10}$ $0.15_{\pm 0.12}$ $3731.50_{\pm 5.08}$ 90%
Table 6: Square Stick Insertion in Real-world Environment. Each entry is evaluated with 20 rollouts in the mean $\pm$ standard deviation format. Falcon is set with $T_{p}=32,T_{a}=16,T_{o}=1,\epsilon=0.02,|\mathcal{B}|=20$.
4.3 Falcon is a Plug & Play Acceleration Algorithm
@@
In the first experiment, we examine our algorithm on the Robomimic benchmarks. We present the success rates, NFE, and speedup for both the original diffusion models, their Falcon-enhanced counterparts, and SDP in Table 1, Table 2, and Table 4, respectively. To better understand Falcon’s impact, we categorize the tasks into simpler and more complex ones based on the level of fine-grained manipulation required. In the Robomimic benchmark, Lift and Can are relatively simple tasks, while the remaining tasks, Square, Transport, and ToolHang, are more complex.
@@
Simple tasks.  The Lift and Can tasks involve smooth and predictable movements. In Lift, the robotic arm moves downward to grasp a small cube and then upward to lift it, with minimal variation in action within each phase. Can follows a similar pattern, where the robot picks up a coke can from a bin and places it in a target bin, primarily involving steady translational motion. Both tasks exhibit strong dependencies between consecutive actions, which makes them ideal for Falcon acceleration. As shown in Fig. 13 (left), Falcon starts denoising with low noise, significantly speeding up the process. Table 4 shows Falcon achieving a 7x speedup in Lift and 2-3x in Can, demonstrating its efficiency in structured motion tasks.
@@
Difficult tasks.  Square, Transport, and ToolHang are more complex and involve finer, more precise movements. For example, in the Transport task, two robot arms must work together to transfer a hammer from a closed container on a shelf into a target bin on another shelf. One arm retrieves the hammer from the container, while the other arm first clears the target bin by removing trash. Then, one arm hands over the hammer to the other, which must place it in the target bin. These actions are dynamic and entail larger changes between consecutive movements. As a result, Falcon can’t start denoising actions with a low noise level (see Fig. 13 (right)). Consequently, Falcon’s speedup is not as pronounced as in simpler tasks. However, even in these more challenging scenarios, as shown in Table 4, Falcon still achieves around 2x speedup while maintaining performance close to that of the original models.
@@
4.4 Falcon mitigates performance degradation in reduced-step solvers
To further demonstrate Falcon’s compatibility with existing acceleration techniques, we compare Falcon-enhanced DPMSolver with a reduced-step DPMSolver baseline (denoted as DPMSolver*), where both use the same number of denoising steps. As shown in Table 3, DPMSolver* exhibits significant performance drops in several tasks when the number of steps is aggressively reduced (e.g., $-10.6\%$ in Square_ph, $-35.6\%$ in Square_mh), while Falcon maintains high success rates with the same number of steps. These results highlight Falcon’s ability to recover performance in low-step regimes and demonstrate its practical value as a plug-and-play solution to boost efficiency without sacrificing accuracy.
Figure 7: Ablation analysis on selection mechanism. This figure compares Falcon’s performance with different partial denoised action selection configurations. Restricting Falcon to choose actions with fixed noise levels (K/2, K/5) significantly reduces score, highlighting that the selection mechanism ensures accurate action choices while accelerating denoising and maintaining high performance.
4.5 Real-World Evaluation of Falcon
To validate the practicality of Falcon beyond simulation, we evaluate it in two real-world robotic manipulation tasks: Dexterous Grasping (Fig. 5) and Square Stick Insertion (Fig. 6).
Dexterous Grasping.  The first task involves dexterous grasping to pick up the blue object placed on a table. Falcon is deployed on top of DDPM without any additional retraining. We compare Falcon+DDPM against SDP(DDPM) and DDPM alone. As shown in Table 5, Falcon achieves a 3.07× speedup over DDPM and outperforms SDP in both runtime (0.145s vs. 0.233s) and CPU memory usage (3730MB vs. 3808MB).
High-Precision Insertion.  The second task involves inserting a square stick into a tall cylindrical chip can. This task requires precise 3D alignment; even slight deviations in height or angle result in failure. Falcon is again applied as a plug-in to DDPM. We train both DDPM and SDP using 50 human demonstrations. As shown in Table 6, Falcon matches DDPM in 90% success rate while being 2.86× faster and more memory efficient.
These results confirm that Falcon not only accelerates diffusion policy inference in simulation but also effectively transfers to real-world robotic applications.
@@ 4.6 Falcon retains the ability to express multimodality policy
We further evaluate Falcon in the PushT task under a symmetric configuration, where multimodality can be visualized directly through trajectory diversity. As shown in Fig. 3, DDPM+Falcon produces diverse reaching trajectories, while Consistency Policy exhibits mode collapse, predominantly favoring a single direction. This qualitative difference demonstrates Falcon’s ability to maintain diverse outputs even in symmetric settings, without retraining.
@@ 4.7 Ablation analysis: sensitivity to the threshold $\epsilon$
We explore how the threshold $\epsilon$ influences Falcon’s performance in the Square_ph task. The threshold $\epsilon$ determines the confidence in reusing partial denoised actions toward the desired action $\boldsymbol{a}_{t:t+T_{p}}$ guided by $\nabla_{\mathbf{A}_{t}^{k}}\log p\left(\mathbf{A}_{t}^{k}\mid\mathbf{O}_{t}\right)$. We conducted 50 evaluations using a range of $\epsilon$ values: $\{10^{-4},5\times 10^{-3},8\times 10^{-3},10^{-2},3\times 10^{-2},5\times 10^{-2},8\times 10^{-2},10^{-1}\}$. As shown in Fig. 4 (left), when $\epsilon$ is small, Falcon is highly selective about using partial denoised actions, leading to slow denoising because Falcon frequently samples from the standard normal distribution. This results in no significant improvement. When $\epsilon$ is large, Falcon becomes overly confident in the partial denoised actions, leading to selecting many incorrect actions that do not align with $\boldsymbol{a}_{t+T_{a}:t+T_{p}}$, which causes a sharp drop in Score. The optimal range of $\epsilon$ lies in the middle, where Falcon can achieve speedup without sacrificing task performance, demonstrating that a balanced $\epsilon$ allows Falcon to maintain both efficiency and accuracy.
4.8 Ablation analysis: sensitivity to exploration rate $\delta$
We assess the impact of the exploration rate $\delta$ on Falcon by running 50 evaluations on the Transport_ph task, using $\delta\in\{0.001,0.05,0.0625,0.076,0.1,0.2,0.33,1\}$. Exploration Rate $\delta$ controls the probability that Falcon starts denoising from $\mathcal{N}(\mathbf{0},\mathbf{I})$ rather than the partial denoised action. This exploration mechanism allows Falcon to fill the latent buffer with more partial denoised actions, increasing exploration and preventing over-reliance on the reference action. As shown in Fig. 4 (right), when $\delta$ is set to 0, Falcon primarily aligns with the reference action at each step. Since the reference action is not perfectly aligned with the desired action $\mathbf{A}_{t}$, errors accumulate over time, causing Falcon to generate incorrect actions, resulting in a significant drop in score. As $\delta$ increases, Falcon samples more diverse actions, improving score by reducing the accumulation of errors. However, further increasing $\delta$ results in a decrease in speed, as the model becomes less reliant on the history of denoised actions and starts exploring more random samples from the normal distribution.
@@ 4.9 Ablation analysis: sensitivity to selection mechanism
We evaluate Falcon’s partial denoised action selection mechanism. Falcon adaptively selects partial denoised actions based on $\epsilon$ and the one-step estimation, estimating whether each action can be denoised to $A_{t}$. We compare this with the case where Falcon is fixed to always sample from actions with noise levels of $K/2$ or $K/5$ (represented by the red and green bars), without considering the adaptive selection mechanism. Fig. 7 (left) shows a significant decrease in Score when Falcon is restricted to these fixed actions, demonstrating that not leveraging the selection mechanism compromises performance. This confirms that Falcon’s Selection Mechanism is crucial for ensuring accurate action choices, allowing the model to maintain high performance while accelerating the denoising process.
@@ 5 Related Work
Diffusion policies are widely used for modeling complex behaviors in robotics. 3D Diffusion Policy (Ze et al., 2024b; Ke et al., ) improves visuomotor performance by integrating 3D visual data from sparse point clouds. Large-scale models like Octo (Team et al., 2024) and RDT (Liu et al., 2024) scale diffusion policy parameters to build vision-language-action foundation models. However, their high computational cost limits real-time applications, necessitating acceleration techniques.
To speed up inference, ODE-based sampling methods like DDIM (Song et al., a) and DPMSolver (Lu et al., 2022) reduce denoising steps, while ParaDiGMS (Shih et al., 2024) leverages parallel computing to accelerate sampling on GPUs. Distillation-based methods further improve efficiency by training models for single-step inference. CP (Prasad et al., 2024) builds on the Consistency Trajectory Model (Kim et al., ), enabling a pre-trained diffusion policy to generate actions within a few steps. ManiCM (Lu et al., 2024) extends this approach to 3D robotic tasks. OneDP (Wang et al., 2024) and SDM (Jia et al., 2024) introduce score-based distillation to reduce performance degradation. However, these methods require task-specific training, limiting adaptability. Streaming Diffusion Policy (SDP) (Høeg et al., 2024) uses partial denoised action trajectories, similar to Falcon. However, SDP requires task-specific training and lacks compatibility with general acceleration techniques.
@@ 6 Conclusion
This paper introduces Falcon, the first diffusion policy approach to accelerate action generation in complex visuomotor tasks by leveraging inter-step dependencies in decision-making. Empirical results confirm that Falcon outperforms strong baselines, such as DDIM and DPMSolver, in terms of speed without sacrificing accuracy or expressiveness. Overall, Falcon offers a simple yet effective solution for real-time robotic tasks, providing efficient action generation for complex visuomotor environments. However, one limitation of Falcon is that it does not use a single set of parameters to accelerate all tasks, since different environments may require task-specific parameter tuning. Future work will explore adaptive parameter selection strategies to enhance its generalizability across diverse robotic applications further.
@@
Acknowledgements
This work is sponsored by National Natural Science Foundation of China (62376013). We sincerely thank Tianjie He for his invaluable support in setting up the hardware platform used in our real-world experiments.
Appendix A Experimental Setup
Figure 8: The hardware platform. Left: Task for Dexterous Grasping. Right: Task for inserting a stick into a chip can
A.1 Simulation Environments
Robomimic (Mandlekar et al., 2022) is a large-scale benchmark designed to evaluate robotic manipulation algorithms using human demonstration datasets. Each task provides two types of human demonstrations: proficient human (PH) demonstrations and mixed proficient and nonproficient human (MH) demonstrations. Each environment uses an action prediction horizon $T_{p}=16$ and an action execution horizon $T_{a}=8$, and all tasks use state-based observations. We construct a CNN-based Diffusion Policy (Chi et al., 2023) with DDPM scheduler using 100 denoising steps and the DDIM/DPMSolver scheduler using 16 denoising steps. Models and are taken from the Diffusion Policy repository: https://diffusion-policy.cs.columbia.edu/data/experiments/low_dim/.
Franka Kitchen (Gupta et al., 2020) is designed for evaluating algorithms on long-horizon, multi-stage robotics tasks. The task involves action sequences of dimension 112 and an episode length of 1200, with an action prediction horizon of $T_{p}=16$ and an action execution horizon of $T_{a}=8$. We construct a transformer-based Diffusion Policy using a DDPM scheduler with 100-step discretization.
Push-T is adapted from IBC (Florence et al., 2022) and involves pushing a block in the shape of a T to a fixed target using a circular end effector. The task features randomized initial poses of the block and end-effector, requiring contact-rich, precise motion. We use state-based observations.
MetaWorld (Yu et al., 2020) is an open-source simulation benchmark for meta-reinforcement learning and multi-task learning. It consists of 50 different robotic manipulation tasks categorized into different difficulty levels ranging from simple to very challenging. We follow the setup and model training in the official 3D Diffusion Policy codebase (Ze et al., 2024b): https://github.com/YanjieZe/3D-Diffusion-Policy.
ManiSkill2 (Gu et al., 2023) is a benchmark for robotic manipulation with 20 tasks, 2000+ objects, and 4M+ demonstrations. It supports rigid/soft-body tasks and fast visual learning (2000 FPS).
A.2 Real-world Experiments
We conduct two real-world manipulation tasks—dexterous grasping and high-precision insertion. As shown in Figure 8, our robot consists of a 7-DoF RealMan RM75-6F arm and a 6-DoF PsiBot hand. It is equipped with a wrist-mounted RealSense D405C camera and a head-mounted RealSense D435 camera. Objects are placed on a table in front of the robot for manipulation. For each real-world task, we collect 50 human demonstrations and use the open-source Diffusion Policy implementation https://github.com/real-stanford/diffusion_policy and SDP implementations https://github.com/Streaming-Diffusion-Policy/streaming_diffusion_policy.
Appendix B Pseudocode of Falcon with other sampling solvers
In this section, we provide the pseudocode for Falcon when integrated with alternative sampling solvers, such as DDIM (Song et al., a) and DPMSolver (Lu et al., 2022).
B.1 Falcon with DDIM
@@
B.2 Falcon with DPMSolver
@@ Appendix C Detail performance of Falcon with 3D Diffusion Policy
Tables 7, 8 and 9 present the success rates, NFE and speedup respectively, for both 3D Diffusion Policy and 3D FalconDDIM. These results further validate Falcon’s effectiveness in reducing inference time while maintaining task performance across different robotic manipulation tasks.